OpenStudy (anonymous):

hey guys, some help please.. if f(x,y) = (x^3+y^4) then f(xy)-f(yx) at the point (1,2) is.. how would i solve this question. Thanks in advance:D

OpenStudy (ayee_ciera):

you would substitute 1 in for x and 2 in for y

OpenStudy (ayee_ciera):

(1^3+2^4)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

is it \[\huge f_{xy}-f_{yx}\]?

OpenStudy (misty1212):

partial derivatives?

OpenStudy (misty1212):

no that can't be right, both second partials are zero must be something else

OpenStudy (anonymous):

lol yes it is. sorry

OpenStudy (misty1212):

both partials are zero everywhere

OpenStudy (misty1212):

\[f(x,y) = (x^3+y^4) \]\] \[f_x(x,y)=3x^2\] so \[f_{xy}(x,y)=0\]

OpenStudy (anonymous):

lol now i'm confused..

OpenStudy (misty1212):

the partial wrt x, treat y as a constant so it is just \(3x^2\) then if you take the partial wrt y, there is no y in it, so it is just zero

OpenStudy (anonymous):

ooh okay sorry about that. thanks:D

OpenStudy (misty1212):

maybe i am not understanding the question correctly

OpenStudy (irishboy123):

is the answer 8?

OpenStudy (anonymous):

its a multiple choice question. i'm not really sure how to approach the question. could it be that the question is incorrect?

OpenStudy (misty1212):

can you post a screen shot or sommat?

OpenStudy (anonymous):

okay just give me a second

OpenStudy (anonymous):

my pc is a little bit slow this evening, please bear with me

OpenStudy (anonymous):

are you able to make that out?

OpenStudy (misty1212):

lol you forgot the exponent!!

OpenStudy (misty1212):

\[f(x,y) = (x^3+y^4)^5\]

OpenStudy (anonymous):

LOL oh right i did, hehe SO SORRY!!

OpenStudy (anonymous):

thats correct:D

OpenStudy (anonymous):

How would i approach this question?

OpenStudy (misty1212):

i think it is still has to be zero, but at now we can compute it first \[f_x=5(x^3+y^3)^4\times 3x^2\] via the chain rule

OpenStudy (misty1212):

is that part clear?

OpenStudy (misty1212):

oops gotta run, you will get more help i am sure

OpenStudy (anonymous):

yeah sort of..

OpenStudy (anonymous):

okay cool thanks:D

OpenStudy (irishboy123):

what exactly are you learning? mixed partials should be the same ordinarily

OpenStudy (anonymous):

partial derivatives

OpenStudy (anonymous):

its a new section of work so im really not too familiar with it

OpenStudy (irishboy123):

ah! so you need to go through the process of doing the derivative in that case, go for it.

OpenStudy (anonymous):

okay cool, thanks dude:D